OpenStudy (anonymous):

Compute the first five terms of the sequence defined recursively by a1=1, a2=3; ak=(k-1)/(k-2). k≥3. The first term is 1, the second is 3.

OpenStudy (anonymous):

this is actually not a recursion the way it is written. it says \[a_k=\frac{k-1}{k-2}\] unless there is something missing

OpenStudy (anonymous):

Oh I see. It's actually: ak=(ak−1)/(ak−2)

OpenStudy (anonymous):

ooh ok

OpenStudy (anonymous):

\[_k=\frac{a_{k-1}}{a_{k-2}}\]

OpenStudy (anonymous):

yes!

OpenStudy (anonymous):

\[a_1=1,a_2=3\] then \[a_3=\frac{a_2}{a_1}=\frac{3}{1}=3\]

OpenStudy (anonymous):

and \[a_4=\frac{a_3}{a_2}=\frac{3}{3}=1\]

OpenStudy (anonymous):

and finally \[a_5=\frac{a_4}{a_3}=\frac{1}{3}=\frac{1}{3}\]

OpenStudy (anonymous):

Oh I see what you did. That makes sense :) Thank you so much! Would you mind helping me with another? Compute the first five terms of the sequence defined recursively by A1=1; An=((2An-1)+2)/(-An-1). n≥2.

OpenStudy (anonymous):

let me try to write it

OpenStudy (anonymous):

\[An=\frac{2A_{n-1}+2}{-A_{n-1}}\]

OpenStudy (anonymous):

like that?

OpenStudy (anonymous):

first replace A by 1 to see what the next term will be. it will be \[A-2=\frac{2\times 1+2}{-1}=-4\]

OpenStudy (anonymous):

i meant \[A_2=-4\]

OpenStudy (anonymous):

Exactly like that!

OpenStudy (anonymous):

the replace A by -4 to find \[A_3=\frac{2\times -4+2}{-(-4)}=-\frac{3}{2}\]

OpenStudy (anonymous):

and so on

OpenStudy (anonymous):

Perfect I see exactly what you're doing. Thank you so much. :) You're great!!